# Hi-EF: Benchmarking Emotion Forecasting in Human-interaction

### v1

@@ Abstract
Affective Forecasting, a research direction in psychology that predicts individuals’ future emotions, is often constrained by numerous external factors like social influence and temporal distance. To address this, we transform Affective Forecasting into a Deep Learning problem by designing an Emotion Forecasting paradigm based on two-party interactions. We propose a novel Emotion Forecasting (EF) task grounded in the theory that an individual’s emotions are easily influenced by the emotions or other information conveyed during interactions with another person. To tackle this task, we have developed a specialized dataset, Human-interaction-based Emotion Forecasting (Hi-EF), which contains 3069 two-party Multilayered-Contextual Interaction Samples (MCIS) with abundant affective-relevant labels and three modalities. Hi-EF not only demonstrates the feasibility of the EF task but also highlights its potential. Additionally, we propose a methodology that establishes a foundational and referential baseline model for the EF task and extensive experiments are provided. The dataset and code is available HERE. 33footnotetext: Engineering Research Center of Robotics, Ministry of Education, Academy for Engineering & Technology, Fudan University, Shanghai, China. 44footnotetext: School of Information Science and Technology, Fudan University, Shanghai, China.
1 Introduction
@@
Affective computing is a branch of computer science and artificial intelligence that aims to enable computers to recognize, interpret, process, and simulate human emotions. Proposed by Rosalind Picard [1], this field combines psychology, computer science, and cognitive science to imbue machines with emotional intelligence. Current research focuses primarily on calculating and analyzing an individual’s present emotional state using facial expressions, voice, and physiological signals, which is Emotion Recognition (ER) [2, 3, 4, 5, 6, 7, 8]. However, despite significant progress in understanding and responding to current emotions, affective computing often overlooks the prediction of future emotions. Affective Forecasting [9] refers to the process by which individuals predict their future emotional reactions to various events. This cognitive mechanism plays a crucial role in decision-making, influencing choices by anticipating potential emotional outcomes. However, empirical studies [9, 10, 11] have shown that individuals often exhibit significant inaccuracies in their affective forecasts. For instance, Gilbert and Wilson [10] highlight the prevalence of “impact bias”, where people tend to overestimate the intensity and duration of their emotional responses to future events. This bias is evident in both positive and negative contexts, such as overestimating happiness from future achievements or distress from anticipated setbacks. Moreover, factors such as focalism, where individuals focus too narrowly on a single event, and underestimating emotional adaptation, contribute to these forecasting errors [11]. Understanding these predictive inaccuracies is critical, as they can lead to suboptimal decision-making in personal, professional, and social contexts. Recent advancements in this field aim to mitigate these biases by exploring the underlying cognitive and emotional mechanisms that drive affective forecasting [9]. Additionally, affective forecasting is influenced by various external factors, one of the most significant being time discounting [12, 13, 14, 15]. Time discounting refers to the tendency of individuals to devalue future events compared to immediate ones. This cognitive bias leads people to underappreciate the emotional impact of distant future events while overestimating the effect of near-term occurrences. As Loewenstein and Prelec [13] describe, individuals exhibit anomalies in intertemporal choice, leading to distorted predictions of future emotional states. Understanding the interplay between affective forecasting and time discounting is crucial for improving decision-making processes, as it helps explain why people often misjudge the long-term emotional consequences of their actions [10].
@@
Research on affective forecasting primarily focuses on psychology and economics. However, we aim to redefine the paradigm of affective forecasting to transform this problem into a Deep Learning challenge. We propose a new task: Emotion Forecasting (EF), predicting one party’s future emotion during a two-party interaction by using short-term contextual information and the current emotional state of the other party. First, emotional exchange is bidirectional: one person’s emotions can affect the other’s, allowing for more accurate predictions of emotional changes [16]. Second, two-party interactions occur over short time frames, making the emotional predictions more reliable as responses to recent events are more direct and intense [13]. Lastly, the dynamic context of interactions provides rich background information, such as conversation content, tone, and emotional shifts, enhancing the precision of affective forecasting [17]. By remodelling affective forecasting to focus on interpersonal emotional dynamics, we can better understand and predict emotional changes in relationships, providing new tools and methods for emotion management and support [11, 10]. And we illustrate the difference between ER task and EF task in Figure 1.
Based on these, we list these criteria to construct dataset for solving EF task: (1) Multi-modal information: By analyzing facial expressions, conversational content, and vocal intonations, we can better assess a person’s emotions, thereby enabling more accurate predictions of the emotions the other party may experience. (2) Strong-Interactive Scenes: Videos should be sourced from highly interactive scenarios, enabling enough amount of interaction data for the dataset. (3) Multilayered-Contextual Information: The data format needs to include three layers of contextual information: contextual information, one party’s current emotional state, and the emotion of the other party that needs to be forecasted.
In adherence to these guidelines, we have developed the Human-interaction-based Emotion Forecasting (Hi-EF). Hi-EF boasts over 3,069 Multilayered-Contextual Interaction Samples (MCIS), featuring in excess of 5,242 video clips with abundant affective-relevant labels. To meet the stringent criteria for data collection, we have devised a comprehensive two-stage strategy to acquire a sufficient pool of candidate MCIS. Furthermore, we have implemented a reliable three-round annotation process to ensure high-quality labeling, given the complexity of the emotional information encapsulated in the labels to make sure the emotion is annotated with high quality. For the challenging EF task, we proposed the baseline method and make comprehensive experiments to demonstrate the feasibility of EF task. Besides that, more experiments of other affective computing tasks are also provided.
@@
Our work makes the following three significant contributions:
  • We have transformed Affective Forecasting from a psychological and economic problem into a Deep Learning task and developed an Emotion Forecasting paradigm based on two-party interactions. This introduces a novel task in the field of affective computing: Emotion Forecasting (EF). Previously, affective computing primarily focused on recognizing current emotions. Our work provides a new perspective for affective by aiming to predict future emotions.
  • To address this novel EF task, we have constructed a Human-interaction-based Emotion Forecasting dataset (Hi-EF). This dataset introduces a unique data format, Multilayered-Contextual Interaction Samples (MCIS), specifically designed for the EF task. MCIS encompasses short-term contextual information, one party’s current emotional state, and the other party’s future emotional state. It includes three modalities and rich affective-relevant labels for both parties. Additionally, our dataset can also be utilized for other affective computing tasks.
  • To generate high-quality MCIS, we developed a sophisticated construction procedure including candidate MCIS generation, anomaly detection and multifaceted-assisted reliable annotation.
  • We have designed a paradigm for the EF task and, based on this paradigm, established a baseline for the EF task. Extensive experiments and analyses were conducted on the Hi-EF dataset, demonstrating the feasibility of the EF task and the significance of the Hi-EF dataset. We will release our method’s code and the dataset for further research and validation.
2 Significance of EF task
@@
  • Individual Emotion Modeling. By analyzing extensive data from a person’s interactions, we can model their emotional responses, identifying what types of interactions are likely to elicit specific emotions. This modeling is particularly useful in the medical field for treating conditions such as depression, enabling personalized therapeutic approaches.
@@
3 Relevant Datasets of Emotion Forecasting: Emotion Recognition Datasets
@@
Text Description Affective Component
ID Datasets Class Scene Fac. Scene Modality Inte. Stre. Emo. Uncer. DFER MER ERC EF
1 IEMOCAP (2008) [18] 5 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
2 HOW (2011) [19] 3 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
3 CMU-MOSEI (2018) [20] 6 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
4 MELD (2018) [21] 7 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ $\checkmark$ -
5 CAER (2019) [22] 7 1 - - V. - - $\checkmark$ - $\checkmark$ - - -
6 DFEW (2020) [23] 7 1 - - V. - - $\checkmark$ - $\checkmark$ - - -
7 FERV39K (2022) [24] 7 22 - - V. - - $\checkmark$ - $\checkmark$ - - -
8 MAFW (2022) [8] 11 1 $\checkmark$ $\checkmark$ V. &A. &T. - - $\checkmark$ - $\checkmark$ $\checkmark$ - -
Video-driven Emotion Recognition datasets Datasets. Video-driven Emotion Recognition datasets are dedicated to identifying human emotions from dynamic visual information. These datasets includes DFER datasets, typically originating from controlled laboratory settings, as exemplified by the CK+ [25] and Oulu-CASIA [26] datasets. Additionally, there are DFER datasets in the wild, such as AFEW [27], Aff-wild [28], AFEW-VA [29], CAER [30], DFEW [23], FERV39k [24]. Datasets like MAFW [8], which offer diverse modalities and emotion labels while still is video-driven, since the annotation is mainly based on the visual modality, while it could also be used in MER task.
Multi-modal Emotion Recognition datasets Datasets. MER datasets focuses not just the facial expression. Instead, they pay more attention to fuse the different modalities to analyze the whole emotional states of a person. And the classes is not only the expressions, but also sentiment including positive, neutral and negative and Activation-Valence-Dominance like IEMOCAP [18], HOW [19], CMU-MOSEI[20] and MELD [21]. Especially for MELD, it is text-based Emotion Recognition dataset which is designed for ERC.
The EF task differs from the ER task, although it does involve elements of ER. In the EF task, we first need to recognize one party’s emotions and then, using the contextual information, predict the potential emotions of the other party. This task structure results in the Hi-EF dataset’s MCIS format being entirely different from typical ER formats. For example, in DFER tasks, the dataset usually consists of single videos labeled with the expressions of the person in the video, typically using a seven-category classification for expressions. In contrast, MCIS includes information from both parties’ interactions. It comprises four consecutive videos: the first two videos provide contextual information, the third video presents one party’s emotional state, and the fourth video is where we need to predict the other party’s emotional state. More details about the EF task, Hi-EF, and MCIS will be elaborated in Section 4.
4 The Hi-EF Dataset
4.1 Paradigm of Emotion Forecasting
@@
  • Contextual Interaction Fusion Function $I$ = $f_{CI}(\cdot)$: This function merges various layers of contextual information, providing a refined backdrop for emotion interpretation and forecast. $I$ indicates contextual information, such as non-verbal cues and the engagement levels of the participants
  • Current Party Emotion Recognition Function $E_{A}$ = $f_{ER}(\cdot)$: It assesses the current emotional state of the Party A, multiple modal information cues. $E_{A}$ is the current emotional state of Party A.
@@
It is important to note that in our subsequent model architecture, we do not explicitly indicate the aforementioned three functions. In our model, these functions are integrated rather than being three independent modules. Nonetheless, the model fundamentally adheres to the paradigm we designed.
4.2 Design of Multilayered-Contextual Interaction Sample (MCIS)
@@
To adapt to the EF task, we needed to define a data format. In the introduction, we also outlined our requirements for the dataset: Multi-modal, Strong-Interactive Scene, and Multilayered-Contextual information. Therefore, we designed the Multilayered-Contextual Interaction Sample (MCIS) as the data format for the Hi-EF dataset. Figure 2 illustrates MCIS. An MCIS primarily consists of the following components: Contextual Information: Video clip I and II; Current Information of Party A: Video clip III; Forecasted Information of Party B: Video clip IV. We provide three modalities for each video clip: video, text (utterance), and audio. Additionally, we provide annotation information for clips III and IV. It is important to note that for the Contextual Information part, we only use two clips. We did not choose more than two videos for contextual information mainly due to the following considerations: 1) In Affective Forecasting, as mentioned [12, 13, 14, 15], the time discounting issue suggests that overly long contextual information might introduce noise, affecting emotion prediction. 2) Excessively long contextual information significantly increases computational costs. 3) Contextual information serves as supplementary data, while the primary data for EF is the current emotional state of Party A [13]. This is supported by our subsequent experiments. We provide the visualization of MCIS in the Appendix.
@@
Figure 3: Overview of the Construction Procedure for the Hi-EF Dataset. a) Shows the process of generating candidate MCIS. b) Illustrates the anomaly detection and the rules and multifaceted-assisted reliable annotation for candidate MCIS. c) Details the annotation process. All other annotation information is intended to support the more accurate labeling of emotions. Naturally, we will provide all the annotation details within the dataset.
4.3 Dataset Construction Procedure.
@@
In this section, we present a comprehensive overview of our database creation process, which unfolds in three distinct stages: the collection of raw data and the generation of candidate video clips, the creation of candidate MCIS, and the subsequent annotation of MCIS.
@@
Data Collection and Candidate MCIS Generation. Our dataset is sourced raw video from TV shows, rich in interactive scenes. We then extract subtitles to segment videos into clips based on the timestamp of the utterances. However, these initial clips required refinement to fit MCIS due to excess noise. We applied facial detection and a state-of-the-art ER model [31] to isolate emotionally relevant clips for MCIS construction. The purpose of this approach is to obtain video clips rich in emotions. Our methodology involves selecting three preceding clips along with the emotionally significant clip to form an MCIS unit, as detailed in Figure 3 a). And we finally get a large amount of candidate MCIS.
@@
Anomaly Detection Rules. Our candidate MCIS generation strategy, based on utterances, occasionally results in abnormal clips. To counter this, we’ve developed a unique approach for anomaly detection to maintain annotation quality. As detailed in Figure 3 b), We defined three types of abnormalities: 1) Facial Anomalies: Clips lacking frontal facial visibility of the speaker for at least half of their duration. 2) Scene Transitions: MCIS where the final two clips are from different scenes. 3) Interaction Anomalies: MCIS featuring the same individual in the last two clips. We implemented an algorithm for initial screening against the first and third types of abnormalities, effectively identifying and discarding about half of the non-conforming MCIS. Recognizing the limitations of automated detection in our stringent criteria, manual review is essential. Therefore, we combine algorithmic and manual methods for anomaly detection to enhance accuracy, which is also contained in the next step.
@@
Multifaceted-Assisted Reliable Annotation. The Hi-EF database stands apart due to its novel data format and undergoes a detailed three-round manual annotation process, depicted in Figure 3 (b) and (c). This process encompasses the evaluation of video, audio, and text elements. Initially, crowdsourced workers apply the abnormality criteria to sift through candidate MCIS, identifying those that are normal. These normal MCIS are then rigorously reviewed by professional annotators in the second stage. This stage is crucial for validating initial annotations and removing inconsistencies, with annotators receiving feedback to enhance their work. During this round, affective components and textual descriptions are also annotated to support reliable emotion annotation. The final round involves a thorough re-assessment of the double-checked MCIS, focusing on the last two clips of each sequence. Here, the same MCIS is annotated by five different crowdsourcing annotators. A professor, leveraging multifaceted-assisted auxiliary information, meticulously reviews these annotations to select the most appropriate or provide improved annotation. We provide extensive annotation information, including textual descriptions and emotional components. The textual descriptions cover facial actions and scene information, while the emotional components include interaction polarity, interaction intensity, emotion categories, and the degree of annotation uncertainty. These annotations are intended to comprehensively support the accuracy of emotion category labeling and can also be utilized for other tasks. Detailed annotation information will be introduced in the Appendix.
4.4 Dataset Statistics
@@
The Hi-EF dataset has been meticulously compiled from TV shows, amassing over 100 hours of content and dividing it into more than 60,000 video clips. Following preprocessing, 30,000 clips containing emotion have been retained, and through a series of data refinement, a half of the candidates are eliminated. Manual judgment has further refined the dataset, resulting in a selection of 3,069 MCIS and 5,242 video clips with high-quality annotations. More information of dataset statistics are provided in the Appendix.
5 Experiment
5.1 Experiment Setup
@@
Implementation Details. All experiments were implemented using PyTorch on Nvidia RTX 3090 GPUs. We set the learning rates to be in the range of 1e-2 to 1e-5, weight decay to 1e-6, and used a batch size of 8. All models were trained from scratch using optimer (Adam) for 50 epochs, with uniform frame subsampling at an interval of 8 frames. Furthermore, due to the limited availability of data in the context of Hi-EF, we employed several data augmentation techniques, including random cropping, changes in lighting conditions, and image flipping. To reduce the computational demands, all cropped images were resized to 112$\times$112, and the entire images were resized to 112$\times$168.
Evaluation Metrics. Drawing inspiration from prevalent Affective Computing tasks [32, 23, 8, 24, 33, 30, 34], we selected two traditional metrics for evaluation: Weighted Average Recall (WAR) and Unweighted Average Recall (UAR).
@@
Model Architecture. Our model architecture, as depicted in Figure 4, is divided into two main parts: intra-video information fusion and inter-video information fusion. Firstly, intra-video information fusion is further divided into two segments. One segment involves the alignment and fusion of the three modalities: video, audio, and text. The other segment focuses on video modality, where we divide video frames into three parts: facial expressions, body posture, and scene information, as all three contribute to the overall emotional state of a person. For the vision encoder, we use ResNet18, I3D, and ViT. For the audio encoder, we employ the commonly used AudioCLIP Encoder, and for the text encoder, we use the CLIP text encoder. Regarding the intra-video fusion strategy, we adopt two approaches: one is the average method, and the other is fusion using a transformer. For inter-video information fusion, we primarily use average, transformer, and a combination of LSTM and transformer. It is important to note that due to the inability of ResNet18 and I3D to align with the audio encoder and text encoder, we only utilized the video modality for these two experiments.
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ Abstract
Affective Forecasting is an psychology task that involves predicting an individual’s future emotional responses, often hampered by reliance on external factors leading to inaccuracies, and typically remains at a qualitative analysis stage. To address these challenges, we narrows the scope of Affective Forecasting by introducing the concept of Human-interaction-based Emotion Forecasting (EF). This task is set within the context of a two-party interaction, positing that an individual’s emotions are significantly influenced by their interaction partner’s emotional expressions and informational cues. This dynamic provides a structured perspective for exploring the patterns of emotional change, thereby enhancing the feasibility of emotion forecasting.
Introduction
@@
Affective computing is a branch of computer science and artificial intelligence that aims to enable computers to recognize, interpret, process, and simulate human emotions. Proposed by Rosalind Picard (Picard 2000), this field combines psychology, computer science, and cognitive science to imbue machines with emotional intelligence. Current research focuses primarily on calculating and analyzing an individual’s present emotional state using facial expressions, voice, and physiological signals, which is Emotion Recognition (ER) (Lo et al. 2020; Tran et al. 2015; Liu et al. 2020; Ma et al. 2019; Li et al. 2023; Zhang et al. 2022; Liu et al. 2022). However, despite significant progress in understanding and responding to current emotions, affective computing often overlooks the prediction of future emotions. Affective Forecasting (Wilson and Gilbert 2003) refers to the process by which individuals predict their future emotional reactions to various events. This cognitive mechanism plays a crucial role in decision-making, influencing choices by anticipating potential emotional outcomes. However, empirical studies (Wilson and Gilbert 2003; Gilbert and Wilson 2007; Wilson and Gilbert 2005) have shown that individuals often exhibit significant inaccuracies in their affective forecasts. For instance, Gilbert and Wilson (Gilbert and Wilson 2007) highlight the prevalence of “impact bias”, where people tend to overestimate the intensity and duration of their emotional responses to future events. This bias is evident in both positive and negative contexts, such as overestimating happiness from future achievements or distress from anticipated setbacks. Moreover, factors such as focalism, where individuals focus too narrowly on a single event, and underestimating emotional adaptation, contribute to these forecasting errors (Wilson and Gilbert 2005). Understanding these predictive inaccuracies is critical, as they can lead to suboptimal decision-making in personal, professional, and social contexts. Recent advancements in this field aim to mitigate these biases by exploring the underlying cognitive and emotional mechanisms that drive affective forecasting (Wilson and Gilbert 2003).
@@
Research on affective forecasting primarily focuses on psychology and economics. However, we aim to redefine the paradigm of affective forecasting to transform this problem into a deep learning challenge. We propose a new task: Emotion Forecasting (EF), predicting one party’s future emotion during a two-party interaction by using short-term contextual information and the current emotional state of the other party. First, emotional exchange is bidirectional: one person’s emotions can affect the other’s, allowing for more accurate predictions of emotional changes (Hatfield, Cacioppo, and Rapson 1993). Second, two-party interactions occur over short time frames, making the emotional predictions more reliable as responses to recent events are more direct and intense (Loewenstein and Prelec 1992). Lastly, the dynamic context of interactions provides rich background information, such as conversation content, tone, and emotional shifts, enhancing the precision of affective forecasting (Barsade 2002). By remodelling affective forecasting to focus on interpersonal emotional dynamics, we can better understand and predict emotional changes in relationships, providing new tools and methods for emotion management and support (Wilson and Gilbert 2005; Gilbert and Wilson 2007). And we illustrate the difference between ER task and EF task in Figure 1.
@@
Based on these, we list these criteria to construct datasets for solving the EF task: (1) Multi-modal information: By analyzing facial expressions, conversational content, and vocal intonations, we can better assess a person’s emotions, thereby enabling more accurate predictions of the emotions the other party may experience. (2) Strong-Interactive Scenes: Videos should be sourced from highly interactive scenarios, enabling enough amount of interaction data for the dataset. (3) Multilayered-Contextual Information: The data format needs to include three layers of contextual information: contextual information, one party’s current emotional state, and the emotion of the other party that needs to be forecasted.
In adherence to these guidelines, we have developed the Human-interaction-based Emotion Forecasting (Hi-EF) dataset. Hi-EF boasts 3,069 Multilayered-Contextual Interaction Samples (MCIS), featuring in 5,242 video clips with abundant affect-related labels. To meet the stringent criteria for data collection, we have devised a comprehensive two-stage strategy to acquire a sufficient pool of candidate MCIS. Furthermore, we have implemented a reliable three-round annotation process to ensure high-quality labeling, given the complexity of the emotional information encapsulated in the labels to make sure the emotion is annotated with high quality. For the challenging EF task, we proposed the baseline method and made comprehensive experiments to demonstrate the feasibility of EF task. Besides that, more experiments of other affective computing tasks are also provided.
Our work makes the following significant contributions:
  • We have transformed Affective Forecasting from a psychological and economic problem into a deep learning task and developed an Emotion Forecasting paradigm based on two-party interactions. This introduces a novel task in the field of affective computing: Emotion Forecasting (EF). Previously, affective computing primarily focused on recognizing current emotions. Our work provides a new perspective by aiming to predict future emotions.
  • To address this novel EF task, we have constructed a Human-interaction-based Emotion Forecasting dataset (Hi-EF). This dataset introduces a unique data format, Multilayered-Contextual Interaction Samples (MCIS), specifically designed for the EF task. MCIS encompasses short-term contextual information, one party’s current emotional state, and the other party’s future emotional state. It includes three modalities and rich affect-related labels for both parties. Additionally, our dataset can also be utilized for other affective computing tasks.
  • To generate MCIS, we developed a construction procedure including candidate MCIS generation, anomaly detection and multifaceted-assisted reliable annotation.
  • We have designed a paradigm for the EF task and, based on this paradigm, established a baseline for the EF task. Extensive experiments and analyses were conducted on the Hi-EF dataset, demonstrating the feasibility of the EF task and the significance of the Hi-EF dataset. We will soon release our method’s code and the dataset for further research and validation.
Text Description Affective Component
ID Datasets Class Scene Fac. Scene Modality Polar. Inten. Emo. Uncer. DFER MER ERC EF
1 IEMOCAP (2008) (Busso et al. 2008) 5 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
2 HOW (2011) (Morency, Mihalcea, and Doshi 2011) 3 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
3 CMU-MOSEI (2018) (Zadeh et al. 2018) 6 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ - -
4 MELD (2018) (Poria et al. 2018) 7 1 - - V. &A. &T. - - $\checkmark$ - - $\checkmark$ $\checkmark$ -
5 CAER (2019) (Lee et al. 2019a) 7 1 - - V. - - $\checkmark$ - $\checkmark$ - - -
6 DFEW (2020) (Jiang et al. 2020) 7 1 - - V. - - $\checkmark$ - $\checkmark$ - - -
7 FERV39K (2022) (Wang et al. 2022a) 7 22 - - V. - - $\checkmark$ - $\checkmark$ - - -
8 MAFW (2022) (Liu et al. 2022) 11 1 $\checkmark$ $\checkmark$ V. &A. &T. - - $\checkmark$ - $\checkmark$ $\checkmark$ - -
Significance of EF Task
@@
  • Individual Emotion Modeling. By analyzing extensive data from a person’s interactions, we can model their emotional responses, identifying what types of interactions are likely to elicit specific emotions. This modeling is useful in the medical field for treating conditions such as depression, enabling personalized therapeutic approaches.
@@
Relevant Datasets of Emotion Forecasting: Emotion Recognition Datasets
@@
Video-driven Emotion Recognition Datasets. Video-driven Emotion Recognition datasets are dedicated to identifying human emotions from dynamic visual information. These datasets includes DFER datasets, typically originating from controlled laboratory settings, as exemplified by the CK+ (Lucey et al. 2010) and Oulu-CASIA (Zhao et al. 2011) datasets. Additionally, there are DFER datasets in the wild, such as AFEW (Dhall et al. 2018), Aff-wild (Kollias et al. 2019), AFEW-VA (Kossaifi et al. 2017), CAER (Lee et al. 2019b), DFEW (Jiang et al. 2020), FERV39k (Wang et al. 2022a). Datasets like MAFW (Liu et al. 2022), which offer diverse modalities and emotion labels while is still video-driven, since the annotation is mainly based on the visual modality, while it could also be used in MER task.
Multi-modal Emotion Recognition Datasets. MER datasets focus not just the facial expression. Instead, they pay more attention to fuse the different modalities to analyze the whole emotional state of a person. And the class is not only the expression, but also sentiment including positive, neutral and negative and Activation-Valence-Dominance like IEMOCAP (Busso et al. 2008), HOW (Morency, Mihalcea, and Doshi 2011), CMU-MOSEI(Zadeh et al. 2018) and MELD (Poria et al. 2018). Especially for MELD, it is text-based Emotion Recognition dataset designed for ERC.
The EF task differs from the ER task, although it does involve elements of ER. In the EF task, we first need to recognize one party’s emotions and then, using the contextual information, predict the potential emotions of the other party. This task structure results in the Hi-EF dataset’s MCIS format being entirely different from typical ER formats. For example, in DFER tasks, the dataset usually consists of single videos labeled with the expressions of the person in the video, typically using a seven-category classification for expressions. In contrast, MCIS includes information from both parties’ interactions. It comprises four consecutive videos: the first two videos provide contextual information, the third video presents one party’s emotional state, and the fourth video is where we need to predict the other party’s emotional state. More details about the EF task, Hi-EF, and MCIS will be elaborated in Section The Hi-EF Dataset.
The Hi-EF Dataset
Paradigm of Emotion Forecasting
@@
  • Contextual Interaction Fusion Function $I$ = $f_{CI}(\cdot)$: This function merges various layers of contextual information, providing a refined backdrop for emotion interpretation and forecasting. $I$ indicates contextual information, such as non-verbal cues and the level of participation of the participants.
  • Current Party Emotion Recognition Function $E_{A}=f_{ER}(\cdot)$: It recognizes Party A’s current emotional state from multi-modal cues.
@@
It is worth noting that in our model architecture, the three functions are not explicitly defined as independent modules but are integrated into the overall design, which still follows the proposed paradigm.
Design of Multilayered-Contextual Interaction Sample (MCIS)
@@
To adapt to the EF task, we needed to define a data format. In the introduction, we also outlined our requirements for the dataset: Multi-modal, Strong-Interactive Scene, and Multilayered-Contextual information. Therefore, we designed the Multilayered-Contextual Interaction Sample (MCIS) as the data format for the Hi-EF dataset. Figure 2 illustrates MCIS. An MCIS primarily consists of the following components: Contextual Information: Video clip I and II; Current Information of Party A: Video clip III; Forecasted Information of Party B: Video clip IV. We provide three modalities for each video clip: video, text (utterance), and audio. Additionally, we provide annotation information for clips III and IV. It is important to note that for the Contextual Information part, we only use two clips. We did not choose more than two videos for contextual information mainly due to the following considerations: 1) In Affective Forecasting (Frederick, Loewenstein, and O’donoghue 2002; Loewenstein and Prelec 1992; Ainslie 1975; Laibson 1997), overly long contextual information might introduce noise, affecting the affective forecast. 2) Excessively long contextual information significantly increases computational costs. 3) Contextual information serves as supplementary data, while the primary data for EF is the current emotional state of Party A (Loewenstein and Prelec 1992).
@@
Figure 3: Overview of the Construction Procedure for the Hi-EF Dataset. a) Shows the process of generating candidate MCIS. b) Illustrates the anomaly detection and the rules and multi-facet assisted reliable annotation for candidate MCIS. c) Details the annotation process. All other annotation information is intended to support the more accurate labeling of emotions. Naturally, we will provide all the annotation details within the dataset.
Dataset Construction Procedure
@@
In this section, we present an overview of our database construction process, which consists of three stages: raw data collection and candidate clip generation, candidate MCIS creation, and MCIS annotation.
@@
Data Collection and Candidate MCIS Generation. We source raw videos from shows rich in interactive scenes and extract subtitles to segment them into clips based on utterance timestamps. Since the initial clips contain noise, we refine them using facial detection and a state-of-the-art ER model (Tao et al. 2024) to isolate emotionally relevant segments for MCIS construction. Each MCIS is formed by combining three preceding clips with the emotionally significant one, as illustrated in Figure 3 a). This process yields a large number of candidate MCIS samples.
@@
Anomaly Detection Rules. Our utterance-based MCIS generation strategy occasionally yields abnormal clips. To address this, we design an anomaly detection mechanism to ensure annotation quality. As illustrated in Figure 3 b), we define three anomaly types: (1) Facial anomalies — clips lacking a frontal face view of the speaker for at least half their duration; (2) Scene transitions — MCIS whose final two clips belong to different scenes; (3) Interaction anomalies — MCIS featuring the same individual in the last two clips. We implement an algorithm to pre-screen the first and third types, filtering out about half of the invalid MCIS. Given the limitations of automatic detection under strict criteria, manual inspection remains indispensable. Therefore, we combine algorithmic and manual detection to enhance overall accuracy, which also aligns with the subsequent annotation step.
@@
Multifaceted-Assisted Reliable Annotation. The Hi-EF database adopts a novel data format and undergoes a rigorous three-round manual annotation process (Figure 3 b–c) involving video, audio, and text modalities. In the first round, crowdsourced workers apply anomaly criteria to select valid MCIS. In the second round, professional annotators validate and refine these samples, ensuring consistency and annotating affective components and textual descriptions. The final round re-assesses the double-checked MCIS, focusing on the last two clips. Each MCIS is labeled by five independent annotators, and a professor reviews these annotations with multifaceted auxiliary information to determine or refine the final label. We provide comprehensive annotation details, including textual descriptions (facial actions, scene information) and emotional components (interaction polarity, intensity, emotion categories, and uncertainty). These annotations collectively support reliable emotion categorization and can facilitate other affective computing tasks. Detailed annotation guidelines are provided in the Appendix.
Dataset Statistics
@@
The Hi-EF dataset has been meticulously compiled from TV shows, amassing over 100 hours of content and dividing it into more than 60,000 video clips. Following preprocessing, 30,000 clips containing emotion have been retained, and through a series of data refinement, a half of the candidates are eliminated. Manual judgment has further refined the dataset, resulting in a selection of 3,069 MCIS and 5,242 video clips with high-quality annotations. Additional dataset statistics are provided in the Appendix
Experiment
Experiment Setup
@@
[…]
Evaluation Metrics. Drawing inspiration from prevalent Affective Computing tasks (Schuller et al. 2010; Jiang et al. 2020; Liu et al. 2022; Wang et al. 2022a, b; Lee et al. 2019b; Zhang et al. 2017), we selected two traditional metrics for evaluation: Weighted Average Recall (WAR) and Unweighted Average Recall (UAR).
@@
Model Architecture. Our model architecture, as depicted in Figure 4, is divided into two main parts: intra-video fusion block and inter-video fusion block. Firstly, intra-video fusion block is further divided into two segments. One clip involves the alignment and fusion of the three modalities: video, audio, and text. The other clips focuses on the video modality, where we divide video frames into three parts: facial expressions, body posture, and scene information, as all three contribute to the overall emotional state of a person. For the vision encoder, we use ResNet18(He et al. 2016), I3D(Carreira et al. 2019), and ViT(Dosovitskiy et al. 2020). For the audio encoder, we employ the commonly used AudioCLIP(Guzhov et al. 2022) Encoder, and for the text encoder, we use the CLIP(Radford et al. 2021) text encoder. Regarding the intra-video fusion strategy, we adopt two approaches: one is the average method, and the other is fusion using a transformer. For inter-video fusion block, we primarily use average, transformer, and a combination of LSTM(Graves and Graves 2012) and Transformer(Vaswani 2017).
@@
R18 (He et al. 2016) Average 22.22 17.02
R18 R18 R18 Average Average 23.15 18.58
R18 R18 R18 Transformer Average 32.41 19.75
R18 R18 R18 Average Transformer 28.70 21.09
@@
I3D(Carreira et al. 2019) Average 28.57 21.73
I3D I3D I3D Average Average 29.62 21.16
I3D I3D I3D Transformer Average 29.63 21.19
I3D I3D I3D Average Transformer 29.62 22.99
@@
ViT (Dosovitskiy et al. 2020) ViT ViT Transformer LSTM+Transformer 31.37 22.39
ViT ViT ViT Transformer AudioCLIP LSTM+Transformer 33.34 21.27
ViT ViT ViT Transformer CLIP LSTM+Transformer 31.03 21.80